ACTIVATION OF C-H BONDS: PURE ROTATIONAL SPECTROSCOPY OF HZnCH_{3} (X̃ ^{1}A_{1})

\underline{M. A. FLORY}, A. J. APPONI, and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectrum of HZnCH_{3} has been observed in its ground electronic state (X̃ ^{1}A_{1}) using both direct absorption and Fourier transform microwave techniques in the frequency range 18-516 GHz. Twelve rotational transitions of this symmetric top species were recorded in K-ladders up to K = 7. The molecule was synthesized from Zn(CH_{3})_{2} in an AC discharge and also from Zn (vapor) + CH_{4} with a DC discharge. From measurements of the spectra of various isotopic species (^{66}Zn, ^{67}Zn, ^{68}Zn, ^{13}C, and ^{2}H), an accurate structure has been determined. The H-C-H bond angle was found to be 108.7^{∘}, slightly smaller than that in ZnCH_{3} or CH_{4}. In addition, nuclear spin-rotation (I ^{.} J) interactions with the methyl hydrogen nuclei and electric quadrupole coupling from the ^{67}Zn nucleus were resolved in the FTMW spectrum. From these data, hyperfine parameters have been established. The value of eqQ = -109.125(11) MHz indicates that the bonds to zinc are primarily covalent. Detection of this species, especially via the Zn + CH_{4} pathway, is a good indication of the ability of metal atoms to insert into C-H bonds.